The numbers are in the ratio 4:3. If they differ by 18, find these numbers.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Let the numbers be 4x and 3x

According to the question, the two numbers are differed by 18. Thus,

4x – 3x = 18

=> x = 18

4x = 4×18 = 72

3x = 3×18 = 54

Thus, the two numbers are 72 and 54

Thanks